OpenStudy (jack1):

Help please! find the general solution of the following system of equations: dx/dt = 2y - 4x dy/dt = 2y - (5/2)x

OpenStudy (jack1):

@jim_thompson5910 @radar would love some guidance if someone's free please? @Zarkon

OpenStudy (jack1):

@satellite73 @Luis_Rivera calling in the big guns please...? @asnaseer @Agent_Sniffles

OpenStudy (radar):

Sorry, out of my league.

OpenStudy (jack1):

@katlin95 @SWAG @timo86m @goformit100 ... @genius12 @Compassionate @cshalvey would any of you guys have an idea where to start...?

OpenStudy (swag):

Sorry I do not know :/

OpenStudy (anonymous):

what if you divide all of it by dt and take integral? maye idk lol.

OpenStudy (anonymous):

\[\begin{cases} x'=2y-4x\\ y'=2y-\frac{5}{2}x \end{cases}\] As a matrix equation, you'd write \[\begin{pmatrix}x'\\y'\end{pmatrix}=\begin{pmatrix}-4&2\\-\frac{5}{2}&2\end{pmatrix}\begin{pmatrix}x\\y\end{pmatrix}\] Find the eigenvalues: \[\begin{align*}\begin{vmatrix}-4-\lambda&2\\ -\frac{5}{2}&2-\lambda\end{vmatrix}=0~&\Rightarrow~\lambda^2+2\lambda - 3=0\\ ~&\Rightarrow~ (\lambda+3)(\lambda-1)=0\\ ~&\Rightarrow~\lambda_1=-3, \lambda_2=1 \end{align*}\]

OpenStudy (anonymous):

Now for the eigenvectors. Take \(\lambda_1=-3\). Let \(\vec{b_1}=\begin{pmatrix}b_1\\b_2\end{pmatrix}\) be the associated eigenvector. Now you solve for \(\vec{b_1}\): \[\begin{pmatrix} -4-(-3)&2\\ -\frac{5}{2}&2-(-3) \end{pmatrix}\begin{pmatrix}b_1\\b_2\end{pmatrix}=\begin{pmatrix}0\\0\end{pmatrix}\\ \begin{pmatrix} -1&2\\ -\frac{5}{2}&5 \end{pmatrix}\begin{pmatrix}b_1\\b_2\end{pmatrix}=\begin{pmatrix}0\\0\end{pmatrix}\] So, you get the following equations: \[b_1=2b_2\\ \frac{5}{2}b_1=5b_2~\Rightarrow~b_1=2b_2\] Now you choose any non-zero values of \(b_1\) and \(b_2\) that work. It looks like the simplest one is when \(b_2=1\), making \(b_1=2\). So, the first eigenvector is \(\vec{b_1}=\begin{pmatrix}2\\1\end{pmatrix}\).

OpenStudy (anonymous):

Similarly, take \(\lambda_1=1\). Let \(\vec{b_2}=\begin{pmatrix}b_1\\b_2\end{pmatrix}\) be the associated eigenvector. \[\begin{pmatrix} -4-1&2\\ -\frac{5}{2}&2-1 \end{pmatrix}\begin{pmatrix}b_1\\b_2\end{pmatrix}=\begin{pmatrix}0\\0\end{pmatrix}\\ \begin{pmatrix} -5&2\\ -\frac{5}{2}&1 \end{pmatrix}\begin{pmatrix}b_1\\b_2\end{pmatrix}=\begin{pmatrix}0\\0\end{pmatrix}\] You have \[5b_1=2b_2\\ \frac{5}{2}b_1=b_2\] (Again, they're the same equation.) Let \(b_1=2\), making \(b_2=5\). So, your second eigenvector is \(\vec{b_2}=\begin{pmatrix}2\\5\end{pmatrix}\).

OpenStudy (anonymous):

The solution to the system of ODEs is then \[\large\begin{pmatrix}x\\y\end{pmatrix}=C_1e^{\lambda_1}\vec{b_1}+C_2e^{\lambda_2}\vec{b_2}\]

OpenStudy (jack1):

... holy cr@p ! that is awesome! thank you so much!

OpenStudy (jack1):

look, what you wrote above now makes the lectures seem more understandable, so thanks, I'm going to try the next one solo ok, so just to clarify in my mind: we can only work out the general solution (not the exact solution) to the equation because we dont have initial values, yeah...? @SithsAndGiggles

OpenStudy (anonymous):

Yep. If you were given something like \(\begin{pmatrix}x(0)\\y(0)\end{pmatrix}=\begin{pmatrix}2\\0\end{pmatrix},\) only then could you solve for \(C_1\) and \(C_2\). And you're welcome :)